OpenStudy (anonymous):

Find the area of a square with side length 1 – 3x. A. 1 + 9x2 B. 1 - 9x2 C. 1 - 6x + 9x2 D. 1 - 6x - 9x2

Parth (parthkohli):

The area of a square is the side squared. So you square 1 - 3x. \( \color{Black}{\Rightarrow (1 - 3x)^2}\) \( \color{Black}{\Rightarrow (1 - 3x)(1 - 3x)}\)

OpenStudy (theviper):

\[(1-3x)^2\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow 1(1 - 3x) - 3x(1 - 3x) }\)

OpenStudy (theviper):

Area of a square=side*side

Parth (parthkohli):

\( \color{Black}{\Rightarrow 1 - 3x - 3x + 9x^2}\)

OpenStudy (anonymous):

Area of Square = \[(Side Length)^2\] And Remember: \[(a - b)^2 - a^2 + b^2 - 2ab\] Here, a = 1 and b = 3x Put and solve it...

OpenStudy (theviper):

=\[(1)^2+(-3x)^2-2(1)(3x)\]\[=1+9x^2-6x\] therefore C

Parth (parthkohli):

C again -.-